OpenStudy (anonymous):

3√(2x)+x√(8x)-5√(18x)

OpenStudy (anonymous):

show all work.

OpenStudy (helder_edwin):

\[ \large 3\sqrt{2x}+x\sqrt{8x}-5\sqrt{18x} \]

OpenStudy (anonymous):

-12√2x+2x√2x

OpenStudy (helder_edwin):

first \[ \large \sqrt{8x}=\sqrt{2\cdot4x}=2\sqrt{2x} \]

OpenStudy (helder_edwin):

\[ \large \sqrt{18x}=\sqrt{2\cdot9x}=3\sqrt{2x} \]

OpenStudy (anonymous):

okay what's next?

OpenStudy (helder_edwin):

plug in the main expresion.

OpenStudy (helder_edwin):

\[ \large 3\sqrt{2x}+x\sqrt{8x}-5\sqrt{18x}= \] \[ \large 3\sqrt{2x}+x[2\sqrt{2x}]-5[3\sqrt{2x}]= 3\sqrt{2x}+2x\sqrt{2x}-15\sqrt{2x} \]

OpenStudy (helder_edwin):

now add/substract like terms

OpenStudy (anonymous):

ohh.....okay thanks

OpenStudy (helder_edwin):

u r welcome

OpenStudy (anonymous):

is that it or am i missing something?

OpenStudy (helder_edwin):

what did u get?

OpenStudy (anonymous):

\[12\sqrt{2x}+2x \sqrt{2x}\]

OpenStudy (anonymous):

oh. ha nvm i got it.

OpenStudy (helder_edwin):

\[ \large =-12\sqrt{2x}+2x\sqrt{2x}=(2x-12)\sqrt{2x} \]

OpenStudy (anonymous):

ohhh.

OpenStudy (anonymous):

Which would be my answer. That's the final one without all the work. :p

OpenStudy (anonymous):

lol yeaa